The company with the lowest weighted average cost of capital is: CompanyV Company W CompanyX CompanyY...

  1. The company with the lowest weighted average cost of capital is:

    1. CompanyV

    2. Company W

    3. CompanyX

    4. CompanyY

    5. Company Z
Answer #1

Please provide rating...

WACC = Weight of equity* Cost of equity + weight of debt * post tax cost of debt
i ii=i/(1+i) iii=1-ii iv v vi=iv*(v-v) vii viii=vii*iii+vi*ii
Company Debt/equity Debt equity YTM Tax rate Post tax cost of debt Re WACC
V 0.5385 35.00% 65.00% 8% 30% 5.60% 14.30% 11.25%
W 3 75.00% 25.00% 9% 35% 5.85% 15.20% 8.19%
X 0.5625 36.00% 64.00% 6% 35% 3.90% 12.90% 9.66%
Y 0.4925 33.00% 67.00% 6% 30% 4.20% 13.40% 10.36%
Z 4 80.00% 20.00% 11% 30% 7.70% 15.60% 9.28%
Lowest WACC is of W
Correct answer is Company W